Speedy Auto Repairs uses a job-order costing system. The companys direct materials consist of replacement parts installed inCompute Mr. Wilkes total job cost. 3. If Speedy establishes its selling prices using a markup percentage of 60% of its totalComplete the question by entering your answers in the tabs given below. Required 1 Required 2 Required 3 Compute Mr. WilkesComplete the question by entering your answers in the tabs given below. Required 1 Required 2 Required 3 If Speedy establishe
Answer #1

1.

Predetermined overhead rate = Variable overhead cost per direct labor-hour + Fixed overhead cost/Estimated direct labor hours

= 1 + 693,000/42,000

= 1 + 16.5

= $17.5 per direct labor hour

2.

Overhead applied = Actual direct labor hours used x Predetermined overhead rate

= 10 x 17.5

= $175

Job cost sheet (Mr. Wikes)

Direct material

660

Direct labor

175

Overhead applied

175

Total cost assigned to Mr. Wilkes

$1,010

3.

Markup percentage = 60% of total job cost

Mark up = 1,010 x 60%

= $606

Amount charged to Mr. Wilkes = Total cost assigned to Mr. Wilkes + Mark up

= 1,010 + 606

= $1,616

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks